A ball is dropped from a height of 100m. Find the velocity when the ball will reach the ground?

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Taking downwards direction to be positive direction,
Displacement of body during whole motion=h=-300m
Acceleration of the body=g=-10m/s^2

From the equation of motion,
v^2-u^2=2as
\implies v^2-0^2=2\times -10\times -100
\implies v=20\sqrt{5}m/s